OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals 1/2(4r+1) -r=2/5(2r-3)+2/3 I Need to show all work

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(4r+1)-r=\frac{ 2 }{ 5 }(2r-3)+\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

\[4r+1-2r+3-r=2/5+3/2-1/2\]

OpenStudy (anonymous):

\[2r+1/2-r=4/5r-6/5+3/2\] \[r+1/2=4/5r+3/10\] \[r-4/5r=3/10-1/2\] \[1/5r=-2/10\] \[r=-1\]